OpenStudy (anonymous):

solve the following equation. Express any irrational solutions in simplest form. (x^2+5x-7)(x+2)=0

OpenStudy (unklerhaukus):

you have an expression , that is a product of two terms, and it equals zero, this implies that at least one of the terms Must be zero

OpenStudy (unklerhaukus):

so (x^2+5x-7)=0 or (x+2)=0

OpenStudy (anonymous):

I guess you have to factor them out but I am confused. This cannot be factored right ?

OpenStudy (unklerhaukus):

for the first term you'll have to use the quadratic formula

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\\ ~\\ x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

ok thank you. I appreciate it.